OpenStudy (anonymous):

A box contains four letter tiles with the letters M, A, T, and H. A spinner has five equal sections numbered 1 through 5. What is the probability of drawing a tile labeled T from the box and spinning an odd number on the spinner? A. 3/20 B. 1/3 C. 4/9 D. 17/20

OpenStudy (anonymous):

@SolomonZelman @campbell_st @babyboy2332

OpenStudy (campbell_st):

well look at the probabilities of each event P(T) = 1/4 P(odd) = 3/5 the events are independent, so multiply the probabilities

OpenStudy (anonymous):

so the answer is

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well its 1/4 x 3/5 =

OpenStudy (anonymous):

a 3/20 @campbell_st

OpenStudy (campbell_st):

thats correct